# Exploring Frequency-Inspired Optimization in Transformer for Efficient Single Image Super-Resolution

### v3

@@ Abstract
Transformer-based methods have exhibited remarkable potential in single image super-resolution (SISR) by effectively extracting long-range dependencies. However, most of the current research in this area has prioritized the design of transformer blocks to capture global information, while overlooking the importance of incorporating high-frequency priors, which we believe could be beneficial. In our study, we conducted a series of experiments and found that transformer structures are more adept at capturing low-frequency information, but have limited capacity in constructing high-frequency representations when compared to their convolutional counterparts. Our proposed solution, the cross-refinement adaptive feature modulation transformer (CRAFT), integrates the strengths of both convolutional and transformer structures. It comprises three key components: the high-frequency enhancement residual block (HFERB) for extracting high-frequency information, the shift rectangle window attention block (SRWAB) for capturing global information, and the hybrid fusion block (HFB) for refining the global representation. To tackle the inherent intricacies of transformer structures, we introduce a frequency-guided post-training quantization (PTQ) method aimed at enhancing CRAFT’s efficiency. These strategies incorporate adaptive dual clipping and boundary refinement. To further amplify the versatility of our proposed approach, we extend our PTQ strategy to function as a general quantization method for transformer-based SISR techniques. Our experimental findings showcase CRAFT’s superiority over current state-of-the-art methods, both in full-precision and quantization scenarios. These results underscore the efficacy and universality of our PTQ strategy. The source code will be released.
@@ 1 Introduction
In recent years, deep learning has shown remarkable success in SISR [6, 7, 8, 9], thanks to its powerful expressive capabilities. Specifically, the introduction of convolutional neural networks (CNNs) that employ innovative architectures like residual and dense connections has significantly improved reconstruction accuracy [10, 11]. Moreover, integrating attention mechanisms, which effectively combine spatial and channel information, has further advanced SISR [12, 13, 14, 15]. More recently, the emergence of transformer architectures has been pivotal, demonstrating their superior ability to capture long-range dependencies and setting new benchmarks in the SISR field [3, 16, 17, 18, 19]. However, despite these significant advancements, existing methods in SISR are still lacking. Firstly, while these methods have effectively integrated off-the-shelf advanced network structures, such as designing transformer blocks for capturing global information or complex convolutional structures for extracting discriminative features, they often overlook the impact of frequency information in the input image on the overall performance of different architectures. Therefore, we believe that a comprehensive analysis from the frequency domain could be beneficial. Secondly, the computational demands of these transformer-based SISR methods pose a significant challenge to their practical deployment, particularly in scenarios with limited computational resources.
@@ 2.1 CNN-based SISR
Figure 1: Influence of high-frequency information on the performance of CNN and transformer architectures. In figures (a) and (b), dashed lines correspond to CNN methods, while solid lines represent transformer methods. The first two subfigures in each row demonstrate performance when frequency is dropped in the frequency domain, and the last two subfigures of each row show performance when frequency is dropped in the image domain. In figure (c), dashed arrows indicate operations in the frequency domain, and solid arrows represent operations in the image domain. (a) As the high-frequency drop ratio and window size increase, transformer models exhibit a smaller change in PSNR compared to CNN, suggesting their superiority in capturing low-frequency information. (b) With increasing high-frequency drop ratio and window size, transformer models show a more pronounced change in PSNR compared to CNN, indicating their limited ability to reconstruct high-frequency information from low-frequency.
@@ 3 Frequency-perspective Analysis
In this section, we investigate the influence of frequency components on the performance of both CNN-based and transformer-based SR models. For our study, we select CARN [30], IMDN [31], and EDSR [21] as representative examples of CNN structures. For transformer structures, we examine SwinIR [3], CAT [16], and HAT [36]. We adopt commonly used datasets for analysis. By gradually dropping frequency components, we illustrate the influence of frequency variations on the outcomes of these models.
3.1 High-frequency Dropping Process
In this section, we conduct two types of frequency dropping processes to evaluate the performance of CNNs and transformers, as illustrated in Fig. 1c. We elaborate the details below:
Frequency Dropping in the frequency domain: for a given high-resolution (HR) image $X^{HR}$, we perform a fast Fourier transform (FFT) on it to obtain its frequency spectrum. Subsequently, we flatten this spectrum into a sequence and arrange it in ascending order based on the magnitudes. With a sequence length of $L$, we define a threshold determined by the drop ratio $\gamma$, $0\leq\gamma\leq 1$, located at the magnitude corresponding to the position $\gamma\cdot L$. Frequency components with magnitudes below this threshold are set to zero. Following this, we perform an inverse fast Fourier transform (IFFT) to generate the HR image with dropped frequencies, referred to as $X_{drop}^{HR}(\gamma)$. The process can be formulated as
$X_{drop}^{HR}(\gamma)=\text{IFFT}(\text{Drop}(\left|\text{FFT}(X^{HR})\right|,\gamma)).$ (1)
@@ 3.1 High-frequency Dropping Process
Afterward, we downsample $X_{drop}^{HR}(\gamma)$ using bicubic interpolation to obtain the LR version $X_{drop}^{LR}(\gamma)$ (e.g.×$\times 4$ down-sampling). Finally, we employ CNN-based and transformer-based SR models to generate the super-resolved counterpart $X_{drop}^{SR}(\gamma)$.
Frequency Dropping in the image domain: to further explore the performance differences between CNN-based and transformer-based methods, we investigated the technique of implicitly dropping high-frequency components in image domain. Instead of directly dropping these components, we applied a median filter to the input image. The process involves three steps. Initially, for a given HR image $X^{HR}$, we selected various window sizes $\theta\in\{3,5,7,9,11\}$ to facilitate different levels of high-frequency component reduction, resulting in increased blurriness of the input image as the window size expanded. The implicit frequency dropping process is described by:
$X_{drop}^{HR}(\theta)=Filter(X^{HR},\theta),$ (2)
where $X_{drop}^{HR}(\theta)$ represents the filtered HR image. Subsequently, we downsampled $X_{drop}^{HR}(\theta)$ using bicubic interpolation, following the frequency dropping in the frequency domain technique, to derive the LR counterpart $X_{drop}^{LR}(\theta)$. Finally, we employed CNN-based and transformer-based SR models to generate the super-resolved counterpart $X_{drop}^{SR}(\theta)$.
@@
3.2 Analysis of Frequency Impact
To evaluate the performance difference between CNNs and transformers, we conducted an experiment to assess the attention bias of different structures towards frequency components. We calculate the Peak Signal-to-Noise Ratio (PSNR), $P^{D}(\gamma/\theta)$, between $X_{drop}^{SR}(\gamma/\theta)$ and $X_{drop}^{HR}(\gamma/\theta)$. Here, $\gamma/\theta$ denotes the two different processes of dropping high-frequency components as previously described. The PSNR drop ratio is defined as
$R_{drop}^{D}(\gamma/\theta)=\frac{P(0)-P^{D}(\gamma/\theta)}{P(0)},$ (3)
where $P(0)$ represents the PSNR without frequency dropping, measured between $X^{SR}$ and $X^{HR}$. As shown in Fig. 1a, the figures demonstrate that transformers are less sensitive to the dropping of high-frequency components and are adept at capturing low-frequency information, evidenced by their relatively stable PSNR in the face of high-frequency components loss. Additionally, we observe the same trend with increasing window size, which implies a reduction in high-frequency components, further supporting our conclusions.
Furthermore, we conduct another experiment to evaluate the effectiveness of different structures in reconstructing high-frequency information. Specifically, we calculate the PSNR $P^{E}(\gamma/\theta)$ between $X_{drop}^{SR}(\gamma/\theta)$ and $X^{HR}$. The PSNR drop ratio can be expressed as
$R_{drop}^{E}(\gamma/\theta)=\frac{P^{E}(\gamma/\theta)-P(0)}{P(0)}.$ (4)
In Fig. 1b, we observe that as the proportion of discarded high-frequency information increases, the transformer model experiences a larger PSNR change compared to the CNN model. This indicates its limited ability to reconstruct high-frequency information from low-frequency data. Moreover, when analyzing frequency dropping in the image domain, the results indicate that with larger window sizes, transformers exhibit more significant PSNR fluctuations than CNNs. This reflects their relative weakness in reconstructing high-frequency information from low frequency, suggesting that transformers are less capable of reconstructing high-frequency representations compared to CNNs.
Based on these observations, we conjectured that a system could achieve enhanced capability by harnessing both low and high-frequency information through the integration of transformer and CNN structures. Motivated by this, we propose a method that leverages the strengths of both CNN and transformer. Specifically, we introduce CNN information as a high-frequency prior to aid the transformer in refining the global representation.
@@ 4 Perform backpropagation and update $\{l_{refined},u_{refined}\}^{K}$ with learning rate $\eta$;
The BR strategy is detailed in Algorithm 3. We start with the coarse boundary values obtained from the ADC stage as our initial inputs. These initial boundaries are then transformed into learnable parameters, allowing the model to fine-tune them through training. The quantizer’s discrete nature results in direct differentiation leading to zero gradients, which hinders backpropagation. To address this issue, we employ straight-through estimation (STE), which enables us to approximate gradient calculation for these parameters.
Our strategy diverges from the method outlined by [27], which applies intermediate feature regularization and iterative updates of weights and activations. Instead, we focus our regularization efforts solely on the output, synchronizing the updates of weights and activations. This difference is critically important as it removes the necessity to store intermediate feature representations, greatly reducing the memory usage during training and contributing to more stable training.
@@ 6.2 Implementation Details
During the quantization, we adjusted the batch size to 2 and proceeded with training for 10 epochs. The learning rate for 8-bit quantization was set at $2\times 10^{-4}$ and increased to $2\times 10^{-3}$ for 6-bit and 4-bit quantizations to boost their performance. The smoothing parameter $\beta$ was set to 0.9. Throughout this phase, quantization was applied to all convolutional and linear layers, as well as to matrix multiplication operations. Additionally, we ensured that the input and output precision remained at 8 bits. This quantization stage was performed on a single NVIDIA 3090 GPU, utilizing the PyTorch framework.
@@ 6.4.1 Effect of HFERB and SRWAB
Model HFERB SRWAB HFB Concat PSNR
CRAFTconv $\checkmark$ $\checkmark$ 30.79
CRAFTtranformer $\checkmark$ $\checkmark$ 31.12
CRAFTconcat $\checkmark$ $\checkmark$ $\checkmark$ 30.92
CRAFT $\checkmark$ $\checkmark$ $\checkmark$ 31.18
@@ 6.4.3 Effect of HFB
(a) Frequency dropping in the frequency domain
(b) Frequency dropping in the image domain
(c) Without frequency dropping
Figure 9: Comparison of average log-amplitude spectra across three frequency dropping methods. (a) Utilizing a frequency dropping method in the frequency domain with an 80% drop ratio. (b) Employing frequency dropping in the image domain with a window size of 3. (c) In contrast, no frequency dropping is applied. The results clearly demonstrate that our CRAFT method effectively restores high-frequency components, showcasing its robustness under diverse frequency dropping conditions.
@@ 6.4.4 Effect of the High-Frequency Prior
We conducted several experiments to investigate the effectiveness of high-frequency prior. Firstly, we swapped the input of $\bm{Q}$ and $\bm{K}$, $\bm{V}$ in HFB and treated the output of SRWAB as $\bm{Q}$ and the output of HFERB as $\bm{K}$, $\bm{V}$ to verify whether global features are dominant in restoration and high-frequency features only serve as a prior for refining the global representation. As shown in Table V, compared to the original design, swapping the input leads to a significant drop in performance, with a 0.51dB decrease in PSNR. Furthermore, we also performed an experiment to formulate the model as a cascade structure to verify the effectiveness of the design introducing high-frequency priors. As shown in Table V, the CRAFTcascade structure resulted in a performance drop, with a 0.3dB decrease in PSNR compared to CRAFT. These results demonstrate the effectiveness of high-frequency priors in the CRAFT model.
Additionally, we visualized the frequency domain representations in Fig. 8 to demonstrate that integrating a high-frequency prior enhances detail sharpness. The high-frequency improvements were quantified using Fourier transforms, where $\Phi(\emph{w/}\hskip 3.0ptH)$ and $\Phi(\emph{w/o}\hskip 3.0ptH)$ denote the spectra of models with and without the high-frequency prior, respectively. Specifically, the spectrum of both models was defined as:
@@ 6.4.5 Exploring CRAFT‘s Capability in Restoring High-Frequency Components
Given our findings that transformers have a limited ability to reconstruct high-frequency information, as discussed in Sec. 3.2, we conducted three experiments using log-amplitude spectral analyses to assess CRAFT’s capability in this regard. Specifically, our experiments were carried out on the image img092 from the Urban100 dataset, applying an magnification factor of 4. We set bicubic and SwinIR [3] as our baseline to evaluate CRAFT’s performance. To facilitate clearer comparison, we upscaled the curves in the highlighted local region, displayed in the top-right corner. In Fig. 9a, we initially processed the input image by dropping frequency in the frequency domain with a drop ratio of 80%. The resulting curves distinctly indicate that CRAFT outperforms other methods in restoring high-frequency components, even after a substantial portion of the frequency has been dropped out. In Fig. 9b, we applied the frequency dropping method in the image domain with a window size of 3. The results indicate that CRAFT also outperforms other methods and restores more high-frequency components. Furthermore, Fig. 9c demonstrates that, even without frequency dropping on the input image, CRAFT consistently exhibits superior high-frequency components restoration, reinforcing its exceptional capability in this aspect.

### v4

@@ Abstract
Transformer-based methods have exhibited remarkable potential in single image super-resolution (SISR) by effectively extracting long-range dependencies. However, most of the current research in this area has prioritized the design of transformer blocks to capture global information, while overlooking the importance of incorporating high-frequency priors, which we believe could be beneficial. In our study, we conducted a series of experiments and found that transformer structures are more adept at capturing low-frequency information, but have limited capacity in constructing high-frequency representations when compared to their convolutional counterparts. Our proposed solution, the cross-refinement adaptive feature modulation transformer (CRAFT), integrates the strengths of both convolutional and transformer structures. It comprises three key components: the high-frequency enhancement residual block (HFERB) for extracting high-frequency information, the shift rectangle window attention block (SRWAB) for capturing global information, and the hybrid fusion block (HFB) for refining the global representation. To tackle the inherent intricacies of transformer structures, we introduce a frequency-guided post-training quantization (PTQ) method aimed at enhancing CRAFT’s efficiency. These strategies incorporate adaptive dual clipping and boundary refinement. To further amplify the versatility of our proposed approach, we extend our PTQ strategy to function as a general quantization method for transformer-based SISR techniques. Our experimental findings showcase CRAFT’s superiority over current state-of-the-art methods, both in full-precision and quantization scenarios. These results underscore the efficacy and universality of our PTQ strategy. The source code is available at: https://github.com/AVC2-UESTC/Frequency-Inspired-Optimization-for-EfficientSR.git.
@@ 1 Introduction
In recent years, deep learning has shown remarkable success in SISR [6, 7, 8], thanks to its powerful expressive capabilities. Specifically, the introduction of convolutional neural networks (CNNs) that employ innovative architectures like residual and dense connections has significantly improved reconstruction accuracy [9, 10]. Moreover, integrating attention mechanisms, which effectively combine spatial and channel information, has further advanced SISR [11, 12, 13, 14]. More recently, the emergence of transformer architectures has been pivotal, demonstrating their superior ability to capture long-range dependencies and setting new benchmarks in the SISR field [3, 15, 16, 17, 18]. However, despite these significant advancements, existing methods in SISR are still lacking. Firstly, while these methods have effectively integrated off-the-shelf advanced network structures, such as designing transformer blocks for capturing global information or complex convolutional structures for extracting discriminative features, they often overlook the impact of frequency information in the input image on the overall performance of different architectures. Therefore, we believe that a comprehensive analysis from the frequency domain could be beneficial. Secondly, the computational demands of these transformer-based SISR methods pose a significant challenge to their practical deployment, particularly in scenarios with limited computational resources.
@@ 2.1 CNN-based SISR
Figure 1: Influence of high-frequency information on the performance of CNN and transformer architectures. Dashed and solid lines correspond to CNN and transformer methods, respectively. (a) With an increase in the high-frequency drop ratio, transformer models exhibit a smaller change in PSNR compared to CNN, suggesting their superiority in capturing low-frequency information. (b) As the high-frequency drop ratio increases, transformer models show a more pronounced change in PSNR compared to CNN, indicating their limited ability to reconstruct high-frequency information from low-frequency.
@@ 3 Frequency-perspective Analysis
In this section, we explore the performance of CNNs and transformers from a frequency perspective. To investigate the impact of different frequencies on these architectures, we conducted two sets of experiments using four commonly used benchmarks, as detailed in Sec. 3.1. Additionally, we provide a more physically realistic analysis of the frequency-dropping method in Sec. 3.2.
3.1 Analysis of Frequency Impact
For our study, we select CARN [29], IMDN [30], and EDSR [20] as representative examples of CNN structures. For transformer structures, we examine SwinIR [3], CAT [15], and HAT [35]. The process of dropping frequency components is depicted in Fig. 1c. Given a high-resolution (HR) image $X^{HR}$, we apply a Fast Fourier Transform (FFT) to obtain its frequency spectrum. In the frequency domain, the components are shifted such that the low frequencies are centered. We then calculate the Euclidean distance of each frequency component from the center of the spectrum. These distances are sorted, and based on the frequency discard ratio $\gamma$, where $0\leq\gamma\leq 1$, we determine the number of frequencies to discard. Specifically, the most distant $\gamma\cdot L$ components (where $L$ is the total number of frequency components) are set to zero. Finally, we apply an Inverse Fast Fourier Transform (IFFT) to reconstruct the image with the high-frequency components dropped, referred to as $X_{drop}^{HR}(\gamma)$. The formulation for this process is as follows
$X_{drop}^{HR}(\gamma)=IFFT(Drop(D(FFT(X^{HR})),\gamma)),$ (1)
@@ 3.1 Analysis of Frequency Impact
where $D(\cdot)$ represents the Euclidean distance calculation. Afterward, we downsample $X_{drop}^{HR}(\gamma)$ using bicubic interpolation to obtain the LR version $X_{drop}^{LR}(\gamma)$ (e.g. ×$\times 4$ down-sampling). Finally, we employ CNN-based and transformer-based SR models to generate the super-resolved counterpart $X_{drop}^{SR}(\gamma)$.
To analyze the attention bias of different structures towards frequency components, we compute the peak signal-to-noise ratio (PSNR) $P^{D}(\gamma)$ between $X_{drop}^{SR}(\gamma)$ and $X_{drop}^{HR}$. We then plot the PSNR drop trend to visualize the difference between the two structures. As shown in Fig. 1a, the PSNR drop ratio for each drop ratio is defined as
$R_{drop}^{D}(\gamma)=\frac{P(0)-P^{D}(\gamma)}{P(0)},$ (2)
where $P(0)$ represents the PSNR without dropping, calculated between $X^{SR}$ and $X^{HR}$. The figures illustrate that the transformer model exhibits reduced sensitivity to high-frequency information and excels in capturing low-frequency information, as evidenced by the smaller PSNR change compared to the CNN model as the proportion of discarded high-frequency information increases.
Furthermore, we conduct another experiment to evaluate the effectiveness of different structures in reconstructing high-frequency information. Specifically, we calculate the PSNR $P^{E}(\gamma)$ between $X_{drop}^{SR}(\gamma)$ and $X^{HR}$ and plot the performance drop trend as previously depicted. The PSNR drop ratio for each drop ratio can be expressed as
$R_{drop}^{E}(\gamma)=\frac{P^{E}(\gamma)-P(0)}{P(0)}.$ (3)
From Fig. 1b, we observe that as the proportion of discarded high-frequency information increases, the transformer model experiences a larger PSNR change compared to the CNN model, indicating its limited ability to reconstruct high-frequency information from low-frequency.
Our empirical results reveal that, although transformers demonstrate impressive performance in tasks like super-resolution, they exhibit shortcomings in handling high-frequency information—critical for capturing fine details such as textures and edges. In contrast, CNNs, with their localized convolution operations, excel in processing high-frequency content more effectively. These findings suggest that transformers could benefit from the complementary strengths of CNNs to enhance their ability to recover intricate details.
To address this limitation, we propose a method that synergizes the strengths of both architectures. Specifically, we integrate CNN-generated information as a high-frequency prior to guide the transformer in refining its global representation. By capitalizing on the CNN’s proficiency in capturing high-frequency details and the transformer’s ability to model long-range dependencies, our approach yield improve results, as reported in Section 6.
Figure 2: Effect of frequency dropping in the image domain using a mean filter.
@@
3.2 Real-world Imaging System
In real-world imaging systems, linear shift-invariant models are commonly used. The Point Spread Function (PSF) describes how the system responds to a point source of light, which results in blurring of the image [39, 40]. After this blurring occurs, the signal is sampled by the sensors. This sampling process further influences the final captured image.
To provide a more physically realistic exploration, we conducted additional experiments using a mean filter, a linear operator, to better simulate real-world conditions. Specifically, we applied the mean filter to the input HR image. The process involves three steps. Initially, for a given HR image $X^{HR}$, we selected various window sizes $\theta\in\{3,5,7,9,11\}$ to reduce different levels of high-frequency components, leading to increased blurriness as the window size grows. The frequency dropping process can be expressed as:
$X_{drop}^{HR}(\theta)=Filter(X^{HR},\theta),$ (4)
where $X_{drop}^{HR}(\theta)$ represents the filtered HR image. Subsequently, we downsampled $X_{drop}^{HR}(\theta)$ using bicubic interpolation to derive the LR counterpart $X_{drop}^{LR}(\theta)$. Finally, we employed CNN-based and transformer-based SR models to generate the super-resolved counterpart $X_{drop}^{SR}(\theta)$. The experimental results are presented in Fig. 2. The top row shows the original HR image along with its filtered versions for window sizes $3$, $5$, and $7$. The second row illustrates the corresponding frequency domain representations. The bottom row presents the PSNR drop rate for various models on the Urban100 dataset. As demonstrated by this experiment, we reached the same conclusion as in Sec. 3.1.
@@ 4 Perform backpropagation and update $\{l_{refined},u_{refined}\}^{K}$ with learning rate $\eta$;
The BR strategy is detailed in Algorithm 3. We start with the coarse boundary values obtained from the ADC stage as our initial inputs. These initial boundaries are then transformed into learnable parameters, allowing the model to fine-tune them through training. The quantizer’s discrete nature results in direct differentiation leading to zero gradients, which hinders backpropagation. To address this issue, we employ straight-through estimation (STE) [44], which enables us to approximate gradient calculation for these parameters.
Our strategy diverges from the method outlined by [26], which applies intermediate feature regularization and iterative updates the boundaries of weights and activations. Instead, we focus our regularization efforts solely on the output, synchronizing the updates to the boundaries of weights and activations. This difference is critically important as it removes the necessity to store intermediate feature representations, greatly reducing the memory usage during training and contributing to more stable training.
@@ 6.2 Implementation Details
During the quantization, we adjusted the batch size to 2 and proceeded with training for 10 epochs. The learning rate for 8-bit quantization was set at $2\times 10^{-4}$ and increased to $2\times 10^{-3}$ for 6-bit and 4-bit quantizations to boost their performance. The smoothing parameter $\beta$ was set to 0.9. Throughout this phase, simulated quantization was applied to all convolutional and linear layers, as well as matrix multiplication operations. Additionally, we ensured that the input and output precision remained at 8 bits. This quantization stage was performed on a single NVIDIA 3090 GPU, utilizing the PyTorch framework.
Model HFERB SRWAB HFB Concat PSNR
CRAFTconv $\checkmark$ $\checkmark$ 30.79
CRAFTtranformer $\checkmark$ $\checkmark$ 31.12
CRAFTconcat $\checkmark$ $\checkmark$ $\checkmark$ 30.92
CRAFT $\checkmark$ $\checkmark$ $\checkmark$ 31.18
@@ 6.4.1 Effect of HFERB and SRWAB
6.4.2 Frequency Analysis of HFERB and SRWAB
@@ 6.4.4 Effect of the High-Frequency Prior
We conducted several experiments to investigate the effectiveness of high-frequency prior. Firstly, we swapped the input of $\bm{Q}$ and $\bm{K}$, $\bm{V}$ in HFB and treated the output of SRWAB as $\bm{Q}$ and the output of HFERB as $\bm{K}$, $\bm{V}$ to verify whether global features are dominant in restoration and high-frequency features only serve as a prior for refining the global representation. As shown in Table V, compared to the original design, swapping the input leads to a significant drop in performance, with a 0.51dB decrease in PSNR. Furthermore, we also performed an experiment to formulate the model as a cascade structure to verify the effectiveness of the design introducing high-frequency priors. As shown in Table V, the CRAFTcascade structure resulted in a performance drop, with a 0.3dB decrease in PSNR compared to CRAFT. These results demonstrate the effectiveness of high-frequency priors in the CRAFT model. Additionally, we visualized the frequency domain representations in Fig. 9 to demonstrate that integrating a high-frequency prior enhances detail sharpness. The high-frequency improvements were quantified using Fourier transforms, where $\Phi(\emph{w/}\hskip 3.0ptH)$ and $\Phi(\emph{w/o}\hskip 3.0ptH)$ denote the spectra of models with and without the high-frequency prior, respectively. Specifically, the spectrum of both models was defined as:
@@ 6.4.5 Exploring CRAFT‘s Capability in Restoring High-Frequency Components
Given our findings regarding the limited ability of transformers to reconstruct high-frequency information, as discussed in Sec. 3, we conducted an experiment using log-amplitude spectral analysis to assess CRAFT’s capabilities. Specifically, we performed this experiment on the image img092 from the Urban100 dataset, applying a magnification factor of 4. Bicubic interpolation and SwinIR [3] were set as baselines to evaluate CRAFT’s performance. The experimental results are depicted in Fig. 10. The x-axis represents increasing frequency values, with higher values indicating higher frequencies. For clearer comparison, we have magnified a localized region of the curves, shown in the top-right inset of the figure. As observed in the figure, the HR image retains the entire range of high-frequency components, while the bicubic interpolation loses nearly all high-frequency details. SwinIR [3] is able to restore some high-frequency information, but CRAFT demonstrates superior performance, recovering a greater extent of high-frequency components.
Figure 10: Comparison of average log-amplitude spectra with different methods. The results clearly demonstrate that our CRAFT method effectively restores high-frequency components.